OpenStudy (anonymous):

how do you solve: 3 = √u-3 - √u

OpenStudy (anonymous):

I'm a little confused with the notation, from how I read it, square both sides and get 9 = u - 3 - u.. which would come out as 9=3.. which is false..

OpenStudy (anonymous):

well the square root sign goes over the whole u-3 if that makes sense

OpenStudy (anonymous):

it's loos like complex answer...?

OpenStudy (anonymous):

\[(3+\sqrt{u})=\sqrt{u-3}\] \[9+6\sqrt{u}+u=u-3\] \[6\sqrt{u}=-12\] \[\sqrt{u}=-1/2\]